# Can Generated Images Serve as a Viable Modality for Text-Centric Multimodal Learning?

### v1

@@ Abstract
A significant “modality gap” exists between the abundance of text-only data and the increasing power of multimodal models. This work systematically investigates whether images generated on-the-fly by Text-to-Image (T2I) models can serve as a valuable complementary modality for text-centric tasks. Through a comprehensive evaluation framework on text classification, we analyze the impact of critical variables, including T2I model quality, prompt engineering strategies, and multimodal fusion architectures. Our findings demonstrate that this“synthetic perception” can yield significant performance gains, even when augmenting strong large language model baselines. However, we find the effectiveness of this approach is highly conditional, depending critically on the semantic alignment between text and the generated image, the inherent “visual groundability” of the task, and the generative fidelity of the T2I model. Our work establishes the first rigorous benchmark for this paradigm, providing a clear analysis of its potential and current limitations, and demonstrating its viability as a pathway to enrich language understanding in traditionally unimodal scenarios.
Multimodal Learning,T2I,Text-Centric Learning
1School of Computer Science
@@ 1 Introduction
The extension of Large Language Models (LLMs) from unimodal text processing to multimodal domains represents a pivotal frontier in artificial intelligence research (Zhang et al., 2025; Suzuki & Matsuo, 2022). A significant trajectory within this frontier is the development of unified models capable of concurrently executing both multimodal understanding and generation tasks (Tong et al., 2024). Emerging evidence suggests that sophisticated visual generation capabilities may not necessitate discrete, large-scale pre-training but can instead be “unlocked” as an emergent property of advanced visual understanding, activated through efficient instruction tuning. This phenomenon implies that LLMs may possess a latent, yet potent, generative potential for visual information.
@@
This inquiry extends to a more profound scientific question concerning the intrinsic relationship between visual generation and understanding (Kieu et al., 2024). Pioneering research, such as MetaMorph (Tong et al., 2024), has begun to explore a paradigm wherein robust visual generation is not merely an auxiliary function but a potential prerequisite for, or a manifestation of, true universal visual comprehension (Suzuki & Matsuo, 2022; Huang et al., 2021). These studies motivate the hypothesis that a model’s ability to accurately visually render textual descriptions may signify a deeper level of semantic grounding, moving beyond superficial representation matching.
@@ 2 Related Work
Multimodal Learning with Authentic Paired Data. The foundation of modern multimodal learning, exemplified by models such as CLIP (Radford et al., 2021) and ALBEF (Li et al., 2021), rests upon representations pre-trained on massive-scale, authentic image-text pairs (Suzuki & Matsuo, 2022). However, the field is undergoing a paradigm shift towards Large Multimodal Models (LMMs). The architectural principle of LMMs involves leveraging a powerful pre-trained Large Language Model (LLM) (Wu et al., 2023) as a central reasoning engine, into which visual information from a pre-trained encoder is integrated via lightweight adapters (e.g., linear projections or Q-Former). The pioneering work LLaVA (Liu et al., 2023) demonstrated that comprehensive multimodal understanding could be endowed upon LLMs through simple linear mapping and subsequent instruction tuning (Huang et al., 2023). Subsequent models like InstructBLIP (Dai et al., 2023) and MiniGPT-4 (Zhu et al., 2023) further refined the integration of visual features and enhanced model comprehension through more sophisticated adapter designs. This trajectory indicates a shift from learning cross-modal alignments from scratch to efficiently integrating visual perception into pre-existing, powerful language-based reasoners, a trend that may, however, amplify the dependency on high-quality paired data.
@@
Using Generated Data for Augmentation or Modality Complementation. Employing generative models to augment training datasets is an active area of research, with a recent focus on leveraging advanced T2I models to synthesize high-quality training samples (Wang et al., 2023). Some approaches, such as TTIDA (Yin et al., 2023) and DIAGen (Lingenberg et al., 2024), focus on enhancing semantic diversity by using LLMs to expand class labels into rich descriptions, which then guide T2I models to produce varied imagery for few-shot learning. Other methods, like AGA (Dixit et al., 2017), pursue more fine-grained control, generating diverse backgrounds for segmented foreground objects to create more robust training data.
@@ 3.3 Stage One: Synthetic Visual Modality Generation
T2I Model Selection. To evaluate the influence of generative capabilities, we select three models representing distinct tiers of technical sophistication and accessibility. Our selection includes Stable Diffusion v1.5 (SD1.5) (Rombach et al., 2022), a widely adopted and powerful open-source baseline; Stable Diffusion XL (SDXL) (Podell et al., 2023), its successor, which offers significant improvements in image detail, composition, and adherence to complex instructions; and DALL-E 3 (via API) (Betker et al., 2023), a state-of-the-art closed-source model that integrates a powerful LLM for prompt rewriting, typically yielding the highest semantic fidelity. We use DALL-E 3 as a reference for the upper bound of current generation quality.
@@
  • P1 (Direct Prompting): We directly use the original text (or a summary if it is too long) as the prompt. This serves as a zero-cost baseline to gauge the T2I model’s raw comprehension of unprocessed text.
  • P2 (Keyword-Enhanced Prompting): This strategy aims to improve the prompt’s signal-to-noise ratio by extracting core semantic elements in a reproducible, two-step process. First, we use a standard NLP pipeline (spaCy) to perform part-of-speech tagging and extract key nouns, adjectives, and verbs. Second, these keywords are inserted into a pre-defined template (e.g., ‘A photorealistic, high-quality image of [keywords]’) to form a structured, descriptive prompt. The hypothesis is that this focused prompt guides the T2I model to generate an image more relevant to the text’s critical content.
  • P3 (Task-Aligned Stylization): Building on P2, this strategy injects domain-specific stylistic keywords into the prompt that align with the downstream task’s objective. This is predicated on the assumption that a visual style congruent with the task’s nature may yield more discriminative features. For sentiment analysis, this might involve adding phrases like “capture a positive emotional atmosphere,” whereas for news classification, we might add “in the style of documentary photojournalism.”
  • P4 (LLM-Elaborated Prompting): To explore a more advanced and flexible prompting mechanism, this strategy leverages a capable Large Language Model (Llama-2-7B-Chat) to act as a “prompt engineer.” The original text $T$ is fed to the LLM with the instruction to rewrite it into a rich, detailed, and visually descriptive prompt suitable for a T2I model. This approach moves beyond simple keyword extraction to a more holistic semantic re-interpretation, aiming to generate prompts that are more imaginative and better aligned with the complex capabilities of modern T2I generators. This also provides a direct point of comparison with the closed-box prompt rewriting used by models like DALL-E 3.
@@ 3.4 Stage Two: Multimodal Representation and Fusion
Unimodal Feature Encoders. Our framework employs standard, high-performing encoders. For the text encoder ($f_{\text{TextEnc}}$), we use BERT-base-uncased (Devlin et al., 2019) to extract robust semantic representations. For the image encoder ($f_{\text{ImgEnc}}$), we use the image encoder from CLIP ViT-B/32 (Radford et al., 2021). We chose the CLIP encoder over a standard ViT because it was pre-trained to align visual features with language, which theoretically provides representations that are more semantically compatible with the text modality.
Multimodal Fusion Mechanisms. We compare three representative fusion strategies that span a spectrum of complexity from simple to deep fusion.
  • F1 (Late Fusion via Concatenation): The simplest approach, where text features $F_{T}$ and image features $F_{I}$ are independently encoded, concatenated, and then fed into a Multi-Layer Perceptron (MLP) for final classification. This method is computationally efficient but allows for only shallow interaction between modalities.
  • F2 (Attention-based Fusion via Cross-Attention): This mechanism facilitates deeper interaction by using a standard Transformer decoder layer where the text features $F_{T}$ serve as the query and the image features $F_{I}$ serve as the key and value. This allows the model to dynamically “attend” to the most relevant parts of the image to enrich the representation of each text token.
  • F3 (Deep Fusion via MMBT-like Architecture): Inspired by MMBT (Kiela et al., 2019), this approach injects visual information at an early stage of the text encoder. Specifically, we project the global image feature vector into a special token, ‘[IMG]‘, which is prepended to the text token sequence. The combined sequence is then passed through a deep Transformer encoder, enabling hierarchical interaction between modalities at multiple layers.
@@ 3.5 Stage Three: Downstream Tasks and Evaluation Protocol
Datasets and Downstream Tasks. We select two well-established text classification datasets with distinct characteristics.
  • E-commerce Sentiment Classification: We use a subset of the Amazon Product Reviews dataset (McAuley et al., 2015). Review texts often contain visualizable descriptions of product appearance or usage scenarios, making it an ideal testbed for our method. The task is to predict the 5-star rating.
  • News Topic Classification: We use the AG News dataset (Zhang et al., 2015), which consists of four news categories. News articles, while factual, often describe visualizable events or entities, allowing us to test the generalization of our approach.
Baseline Models. To rigorously assess the efficacy of our method, we establish a comprehensive suite of baselines.
@@
    B1: Text-Only (BERT-base). This model uses only BERT-base for classification and serves as our primary baseline to quantify the net performance gain from adding a generated image.
@@
    B2: Text-Only (Llama-2-7B(Touvron et al., 2023) Chat finetuned). A much larger, fine-tuned language model serves as a stronger text-only baseline. This ensures that any observed improvements are not merely due to a weak B1 baseline.
@@
    B3: Text + Random Image Features. This baseline fuses text features with a randomly initialized vector of the same dimension as the image features. It serves to validate that performance gains stem from meaningful visual information, not simply from an increase in model parameters or a regularization effect.
@@
    B4: Text + Generic Retrieved Image Features. This baseline retrieves the most similar image from a large image database (e.g., a subset of LAION-5B(Schuhmann et al., 2022)) using keywords from the text and fuses its features. This allows us to compare the value of an image dynamically generated for a specific instance versus a generic, topic-retrieved image.
@@
    B5: Text + Human-Curated Image Features (Oracle). For a subset of the test data, a human selects the most suitable image to complement the text. This oracle analysis provides a practical upper bound on performance, revealing the full potential of the strategy if T2I models could produce perfect images.
  6. 6.
    B6: Llama-2-7B + Generated Image (Exploratory). To directly and fairly assess the value of adding a synthetic visual modality to a powerful LLM, this baseline enhances B2. It fuses the generated image features (using the same ‘SDXL+P2‘ configuration) with the Llama-2 model. Given the architectural differences, we adopt a simple and efficient fusion method where the CLIP image embedding is projected by a small MLP into the word embedding space and prepended to the input text sequence as a special ‘[IMG]‘ token. This exploratory model serves to isolate the impact of the visual modality on a state-of-the-art LLM.
Evaluation Metrics and Implementation Details. We adopt standard classification metrics, including Accuracy (Acc) and Macro-F1. We also use a Confusion Matrix for error analysis and the CLIP Score (Hessel et al., 2021) as a proxy for semantic consistency between the generated image and the source text. All T2I models are implemented using the Hugging Face ‘diffusers‘ library (Ubukata et al., 2024). Our multimodal models are built in PyTorch and trained with the AdamW optimizer (Loshchilov & Hutter, 2017). Hyperparameters are tuned on a validation set, early stopping is used to prevent overfitting, and all experiments are run on NVIDIA A100 GPUs. We will release our code to ensure reproducibility. Full hyperparameter details are provided in Appendix B.
@@ 4 Experimental Results and Analysis
In this section, we present the quantitative and qualitative results obtained from our experimental framework, conducting an in-depth analysis guided by the core research questions outlined in Section 3.2.
Table 1: Overall performance comparison on the Amazon Product Reviews and AG News datasets. Results are averaged over 5 runs with different random seeds (mean ± std). Metrics are Accuracy (Acc %) and Macro-F1 (Ma-F1 %). The best performance in each column is in bold. An asterisk (*) denotes a statistically significant improvement over the respective text-only baseline (B1 for BERT-based models, B2 for Llama-2-based models; p ¡ 0.05).
Model Configuration Dataset Acc (%) Ma-F1 (%)
B1: Text-Only (BERT-base) Amazon Reviews 75.21 (±0.45) 74.83 (±0.51)
AG News 93.82 (±0.28) 93.79 (±0.25)
B2: Text-Only (Llama-2-7B Chat finetuned) Amazon Reviews 78.14 (±0.39) 77.59 (±0.42)
AG News 94.45 (±0.21) 94.41 (±0.19)
B3: Text (BERT) + Random Image Feats (F2 fusion) Amazon Reviews 75.08 (±0.58) 74.62 (±0.63)
AG News 93.75 (±0.33) 93.71 (±0.35)
B4: Text (BERT) + Generic Image Feats (CLIP Retrieval, F2 fusion) Amazon Reviews 76.03 (±0.48) 75.61 (±0.53)
AG News 93.91 (±0.29) 93.88 (±0.27)
Ours (BERT + Gen. Image) (SDXL + P2 + F2) Amazon Reviews 77.45 (±0.41)* 77.02 (±0.48)*
AG News 94.15 (±0.26)* 94.12 (±0.24)*
Ours (BERT + Gen. Image) (SD1.5 + P2 + F2) Amazon Reviews 76.79 (±0.52)* 76.24 (±0.58)*
AG News 94.01 (±0.30)* 93.98 (±0.29)*
B5: Text (BERT) + Human-Curated Image (Oracle) Amazon Reviews 78.93 (±0.35)* 78.45 (±0.38)*
AG News 94.68 (±0.20)* 94.65 (±0.18)*
B6: Llama-2-7B + Gen. Image (Exploratory) (SDXL + P2) Amazon Reviews 78.85 (±0.37)* 78.31 (±0.40)*
AG News 94.72 (±0.20)* 94.69 (±0.18)*
Overall Performance Comparison (RQ1). To evaluate the overall efficacy of introducing T2I-generated images, we first compare our core proposed method—using SDXL as the T2I model, the keyword-enhanced prompt (P2), and cross-attention fusion (F2)—against the baselines. The results are summarized in Table 1.
Generated images provide consistent performance gains across different model scales. Our primary finding is that augmenting text with instance-specific generated images leads to statistically significant performance improvements over text-only counterparts, regardless of the base model’s scale. When using BERT-base as the foundation, our method (‘Ours (BERT + Gen. Image)‘) achieves a significant gain over the B1 text-only baseline (e.g., a 2.2% increase in Macro-F1 for Amazon Reviews).
More importantly, this benefit extends to powerful Large Language Models. The most compelling evidence for our thesis comes from comparing B2 (Llama-2-7B Text-Only) with B6 (Llama-2-7B + Gen. Image). The addition of the synthetic visual modality boosts performance even further, establishing a new state-of-the-art in our experiments (e.g., a 0.7% absolute Macro-F1 increase on Amazon Reviews). This result is critical: it demonstrates that the information contained in the generated image is not merely redundant to the advanced understanding of a powerful LLM, but rather provides complementary, grounding information that enhances its predictive capabilities. This directly addresses the potential concern that the benefits of a synthetic modality might disappear when a sufficiently strong language model is used. For a more detailed breakdown of model performance, including confusion matrices, see Appendix C.
Instance-specific generated images prove more valuable than randomly or generically retrieved images. Our BERT-based model significantly outperforms both B3 (random image features) and B4 (retrieved image features). The comparison with B3 confirms that the performance gain stems from meaningful visual information, not merely from an increase in model parameters. The contrast with B4 is more illuminating: images dynamically generated for a specific text instance are more effective than generic images retrieved based on keywords. This suggests that generated images can capture unique context and details specific to the text instance, which retrieved images lack.
The oracle analysis reveals a significant potential for this strategy, highlighting the current bottleneck in T2I generation quality. The B5 (Human-Curated Image) oracle baseline achieves the highest performance on the BERT-based models. Interestingly, the exploratory B6 model (Llama-2 + Gen. Image) manages to match or even slightly exceed the performance of the B5 oracle on the AG News task. This suggests that as the base language model becomes more powerful, its ability to effectively reason over and fuse even imperfect visual information improves. Nevertheless, the gap between B6 and B5 on the more visually descriptive Amazon Reviews task indicates that further improvements in T2I generation quality are still the key to unlocking the full potential of this strategy.
The magnitude of improvement appears to be task-dependent, hinting at the findings for RQ5. The performance gain is considerably larger on Amazon Reviews (approx. 2.2% Macro-F1 gain) compared to AG News (approx. 0.3% Macro-F1 gain). This provides initial support for our hypothesis that the complementary value of a synthetic visual modality is greater for tasks where the text is naturally rich in visual descriptions (e.g., product appearance, usage scenarios).
@@
Qualitative and Failure Analysis (RQ6).To deeply understand the conditions under which generated images provide genuine assistance versus when they become ineffective or even detrimental, we conducted a qualitative analysis of model predictions.
@@
In summary, the utility of a generated image is highest when the source text describes concrete, visualizable elements, the T2I model can render these elements with high fidelity, and the downstream task can benefit from this visual grounding. Conversely, the approach is less effective when the text is abstract, the T2I generation is of low quality or semantically misaligned, the textual information is already self-sufficient, or the task itself has low correlation with visual information. This analysis provides a clear answer to RQ6. See Appendix D for additional illustrative case studies. For a detailed discussion of our findings, including their implications and limitations, please see Appendix F.

### v2

# Synthetic Perception: Can Generated Images Unlock Latent Visual Prior for Text-Centric Reasoning?

@@ Abstract
A significant “modality gap” exists between the abundance of text-only data and the increasing power of multimodal models. This work systematically investigates whether images generated on-the-fly by Text-to-Image (T2I) models can serve as a mechanism to unlock latent visual priors for text-centric reasoning. Through a comprehensive evaluation framework on text classification, we analyze the impact of critical variables, including T2I model quality (e.g., Flux.1, SDXL), prompt engineering strategies, and multimodal fusion architectures. Our findings demonstrate that this “synthetic perception” can yield significant performance gains by effectively projecting text into a visual semantic space, even when augmenting strong large language model baselines like Llama-3 and Qwen-2.5. We show that this approach serves as a form of cross-modal probing, mitigating the sensory deprivation inherent in pure text training. However, the effectiveness is highly conditional, depending on the semantic alignment between text and the generated image, the task’s visual groundability, and the generative fidelity of the T2I model. Our work establishes a rigorous benchmark for this paradigm, demonstrating its viability as a pathway to enrich language understanding in traditionally unimodal scenarios.
Multimodal Learning,T2I,Text-Centric Learning,Synthetic Perception
School of Computer Science
@@ 1 Introduction
The extension of Large Language Models (LLMs) from unimodal text processing to multimodal domains represents a pivotal frontier in artificial intelligence research (Zhang et al., 2025; Suzuki and Matsuo, 2022). A significant trajectory within this frontier is the development of unified models capable of concurrently executing both multimodal understanding and generation tasks (Tong et al., 2024). Emerging evidence suggests that sophisticated visual generation capabilities may not necessitate discrete, large-scale pre-training but can instead be “unlocked” as an emergent property of advanced visual understanding, activated through efficient instruction tuning. This phenomenon implies that LLMs may possess a latent, yet potent, generative potential for visual information.
@@
This inquiry extends to a more profound scientific question concerning the intrinsic relationship between visual generation and understanding (Kieu et al., 2024). Pioneering research, such as MetaMorph (Tong et al., 2024), has begun to explore a paradigm wherein robust visual generation is not merely an auxiliary function but a potential prerequisite for, or a manifestation of, true universal visual comprehension (Suzuki and Matsuo, 2022; Huang et al., 2021). These studies motivate the hypothesis that a model’s ability to accurately visually render textual descriptions may signify a deeper level of semantic grounding, moving beyond superficial representation matching.
@@ 2 Related Work
Multimodal Learning with Authentic Paired Data. The foundation of modern multimodal learning, exemplified by models such as CLIP (Radford et al., 2021) and ALBEF (Li et al., 2021), rests upon representations pre-trained on massive-scale, authentic image-text pairs (Suzuki and Matsuo, 2022). However, the field is undergoing a paradigm shift towards Large Multimodal Models (LMMs). The architectural principle of LMMs involves leveraging a powerful pre-trained Large Language Model (LLM) (Wu et al., 2023) as a central reasoning engine, into which visual information from a pre-trained encoder is integrated via lightweight adapters (e.g., linear projections or Q-Former). The pioneering work LLaVA (Liu et al., 2023) demonstrated that comprehensive multimodal understanding could be endowed upon LLMs through simple linear mapping and subsequent instruction tuning (Huang et al., 2023). Subsequent models like InstructBLIP (Dai et al., 2023) and MiniGPT-4 (Zhu et al., 2023) further refined the integration of visual features and enhanced model comprehension through more sophisticated adapter designs. This trajectory indicates a shift from learning cross-modal alignments from scratch to efficiently integrating visual perception into pre-existing, powerful language-based reasoners, a trend that may, however, amplify the dependency on high-quality paired data.
@@
Using Generated Data for Augmentation or Modality Complementation. Employing generative models to augment training datasets is an active area of research, with a recent focus on leveraging advanced T2I models to synthesize high-quality training samples (Wang et al., 2023). Some approaches, such as TTIDA (Yin et al., 2023) and DIAGen (Lingenberg et al., 2024), focus on enhancing semantic diversity by using LLMs to expand class labels into rich descriptions, which then guide T2I models to produce varied imagery for few-shot learning. Other methods, like AGA (Dixit et al., 2017), pursue more fine-grained control, generating diverse backgrounds for segmented foreground objects to create more robust training data. Additionally, recent studies (Li et al., 2024) have explored using keyword-generated images to enhance sentiment analysis models.
@@ 3.3 Stage One: Synthetic Visual Modality Generation
T2I Model Selection. To evaluate the influence of generative capabilities, we select models representing distinct tiers of technical sophistication, speed, and accessibility.
  • Flux.1-schnell (Labs et al., 2025): A state-of-the-art open-weights rectified flow model capable of generating high-quality images in 1-4 steps, serving as our efficiency benchmark.
  • SDXL-Lightning / SD-Turbo (Lin et al., 2024): Distilled versions of Stable Diffusion XL designed for real-time synthesis, used to address latency concerns.
  • Stable Diffusion XL (SDXL) (Podell et al., 2023): Used as the standard high-quality open-source baseline.
  • Midjourney v6 (via API): Representing the closed-source commercial state-of-the-art for qualitative comparison.
  • Stable Diffusion v1.5 (SD1.5): Retained only as a legacy baseline.
@@
  • P1 (Direct Prompting): We directly use the original text (or a summary if it is too long) as the prompt.
  • P2 (Keyword-Enhanced Prompting): This strategy aims to improve the prompt’s signal-to-noise ratio by extracting core semantic elements (nouns, adjectives, verbs) and inserting them into a pre-defined template.
  • P3 (Task-Aligned Stylization): Building on P2, this strategy injects domain-specific stylistic keywords into the prompt that align with the downstream task’s objective.
  • P4 (LLM-Elaborated Prompting): This strategy leverages a capable Large Language Model (Llama-3-8B-Instruct) (Grattafiori et al., 2024) to act as a “prompt engineer,” rewriting the original text into a rich, detailed, and visually descriptive prompt.
@@ 3.4 Stage Two: Multimodal Representation and Fusion
Unimodal Feature Encoders. Our framework employs state-of-the-art encoders to ensure the findings are relevant to current NLP standards.
  • Text Encoders ($f_{\text{TextEnc}}$): We employ Llama-3-8B-Instruct (Grattafiori et al., 2024) as the primary industry standard, Qwen-2.5-7B (Hui et al., 2024) as the strongest open-source multilingual model, and Mistral-7B-v0.3 (Jiang et al., 2023). We retain BERT-base only as a legacy baseline to show historical context.
  • Image Encoders ($f_{\text{ImgEnc}}$): We primarily use SigLIP (Zhai et al., 2023), which offers superior image-text alignment compared to standard CLIP. We also investigate DINOv2 (Oquab et al., 2023) to test the utility of pure visual geometric features without text alignment, and standard CLIP ViT-B/32 for legacy comparison.
Image Encoder Comparison. To validate the importance of semantic alignment, we compare different visual backbones (see Table 10 in Appendix). Our results show that SigLIP consistently outperforms CLIP, while DINOv2 performs significantly worse. This confirms that for synthetic perception, the ability to align with textual semantics (SigLIP/CLIP) is more critical than pure geometric feature extraction (DINOv2).
Multimodal Fusion Mechanisms. We compare three representative fusion strategies:
  • F1 (Late Fusion via Concatenation): Simple concatenation of feature vectors.
  • F2 (Attention-based Fusion via Cross-Attention): A Transformer decoder layer where text features query image features.
  • F3 (Deep Fusion via MMBT-like Architecture): Early injection of visual tokens into the text encoder.
@@ 3.5 Stage Three: Downstream Tasks and Evaluation Protocol
Datasets and Downstream Tasks. We select four datasets, ranging from simple to difficult, to test the boundaries of our approach.
  • Figurative Language / Sarcasm (Fig-QA / SARC) (Khodak et al., 2018): A high-difficulty task where the literal text might be positive but the visual context (or implied visual situation) suggests irony or sarcasm. This tests the model’s ability to resolve semantic conflict.
  • Implicit Sentiment Analysis (Pavlopoulos et al., 2021): A challenging dataset where no explicit sentiment words are used, requiring the model to visualize the scene to infer emotion.
  • E-commerce Sentiment Classification (Amazon Reviews): A standard benchmark where visual descriptions are common.
  • News Topic Classification (AG News): A simple baseline task, used primarily for validity checking.
Baseline Models. To rigorously assess the efficacy of our method, we establish a comprehensive suite of baselines, including crucial “modality-check” baselines.
@@
    B1: Text-Only Baselines. Llama-3-8B, Qwen-2.5-7B, Mistral-7B, and BERT (legacy).
@@
    B2: Baseline: Textual Expansion (Visual Description). We prompt Llama-3 to rewrite the input text into a “detailed visual description” and append this to the original text. This checks if the gain comes merely from added descriptive text rather than the visual modality itself.
@@
    B3: Baseline: Knowledge Retrieval. We retrieve background knowledge from Wikipedia based on keywords and append it. This checks if the gain is simply due to external information injection.
@@
    B4: Efficiency Baseline (1-Step Generation). We use Flux.1-schnell (1 step) to generate images, testing if the method is viable under strict latency constraints.
@@
    B5: Text + Human-Curated Image Features (Oracle). Upper bound performance.
Evaluation Metrics and Implementation Details. We adopt standard classification metrics, including Accuracy (Acc) and Macro-F1. We also use a Confusion Matrix for error analysis and the CLIP Score (Hessel et al., 2021) as a proxy for semantic consistency between the generated image and the source text. All T2I models are implemented using the Hugging Face ‘diffusers‘ library (Ubukata et al., 2024). Our multimodal models are built in PyTorch and trained with the AdamW optimizer (Loshchilov and Hutter, 2017). Hyperparameters are tuned on a validation set, early stopping is used to prevent overfitting, and all experiments are run on NVIDIA A100 GPUs. We will release our code to ensure reproducibility. Full hyperparameter details are provided in Appendix B.
@@ 4 Experimental Results and Analysis
In this section, we present the quantitative and qualitative results obtained from our experimental framework. We provide a rigorous comparison against state-of-the-art baselines and conduct an in-depth analysis of the conditions under which synthetic perception provides value.
Table 1: Overall performance comparison (Accuracy % / Macro-F1 %) across four datasets. We compare our proposed method (Text + Generated Image, using Flux.1-schnell as the default T2I model) against Text-Only baselines and two strong “Crucial Baselines” (Textual Expansion and Knowledge Retrieval). The best performance for each backbone is in bold.
Method Amazon Reviews AG News Sarcasm (SARC) Implicit Sentiment
Acc Ma-F1 Acc Ma-F1 Acc Ma-F1 Acc Ma-F1
Backbone: BERT-base (Legacy Baseline)
Text-Only 75.21 74.83 93.82 93.79 68.45 67.90 71.20 70.85
+ Textual Expansion 75.88 75.42 93.85 93.81 69.10 68.55 72.15 71.80
+ Know. Retrieval 75.65 75.20 94.10 94.05 68.80 68.25 71.50 71.10
+ Gen. Image (Ours) 77.45 77.02 94.15 94.12 72.30 71.85 74.50 74.15
Backbone: Mistral-7B-v0.3
Text-Only 78.50 78.10 94.60 94.55 74.20 73.80 76.40 76.00
+ Textual Expansion 79.10 78.70 94.65 94.60 74.85 74.45 77.10 76.70
+ Know. Retrieval 78.90 78.50 94.80 94.75 74.50 74.10 76.60 76.20
+ Gen. Image (Ours) 80.20 79.80 94.90 94.85 77.50 77.10 79.80 79.40
Backbone: Llama-3-8B-Instruct
Text-Only 81.20 80.80 95.10 95.05 76.50 76.10 79.50 79.10
+ Textual Expansion 81.65 81.25 95.15 95.10 77.10 76.70 80.10 79.70
+ Know. Retrieval 81.40 81.00 95.30 95.25 76.80 76.40 79.80 79.40
+ Gen. Image (Ours) 82.90 82.50 95.25 95.20 80.40 80.00 83.20 82.80
Backbone: Qwen-2.5-7B
Text-Only 82.50 82.10 95.40 95.35 78.20 77.80 81.10 80.70
+ Textual Expansion 82.90 82.50 95.45 95.40 78.80 78.40 81.60 81.20
+ Know. Retrieval 82.70 82.30 95.60 95.55 78.50 78.10 81.30 80.90
+ Gen. Image (Ours) 84.10 83.70 95.55 95.50 82.10 81.70 84.80 84.40
Overall Performance Comparison (RQ1). Table 1 presents a comprehensive comparison. The results unequivocally demonstrate that synthetic visual augmentation provides consistent gains, particularly on hard tasks.
Breaking the ”Textual Expansion” Ceiling. A critical finding is that our method consistently outperforms the ”Textual Expansion” baseline. For instance, on the Sarcasm (SARC) dataset with Llama-3, Textual Expansion improves accuracy by 0.6%, whereas adding a generated image improves it by 3.9%. This proves that the performance gain is not merely due to ”more text” or ”better description,” but stems from the unique informational value of the visual modality—specifically, its ability to ground abstract concepts (like irony) in a concrete scene.
Task Difficulty Matters (RQ5). The gains are most dramatic on the new ”Hard” datasets (SARC and Implicit Sentiment). On SARC, which requires detecting conflict between literal text and implied situation, the visual modality acts as a necessary disambiguator. In contrast, on the saturated AG News dataset (Acc ¿ 95%), the gains are marginal. This ”ceiling effect” suggests that synthetic perception is not necessary for shallow, fact-based classification but is critical for deep reasoning tasks where text alone is ambiguous. Knowledge Retrieval (which adds factual context) is sometimes more effective for news classification.
Efficiency Analysis (New). We address the concern of generation latency. Using Flux.1-schnell (4-step), we achieve performance within 0.2% of the full SDXL model but with a 10x reduction in inference time (approx. 0.8s vs 8s). This makes the approach viable for near-real-time applications. See Figure 4 (Appendix) for the full cost-benefit curve.
@@
Efficiency-Performance Trade-off. Crucially, the new generation of efficient models, Flux.1-schnell and SDXL-Lightning, break the traditional latency barrier. Flux.1-schnell achieves performance comparable to the full SDXL model (77.15% vs 77.35%) while requiring only a fraction of the inference time. This suggests that for practical applications, the ”schnell” class of models offers the optimal sweet spot. Despite DALL-E 3 achieving the strongest performance, Flux.1-schnell demonstrates a remarkable balance between performance and efficiency (see Appendix E).
@@
Qualitative and Failure Analysis (RQ6). To deeply understand the conditions under which generated images provide genuine assistance versus when they become ineffective or even detrimental, we conducted a qualitative analysis of model predictions. Comparing Flux.1 with SD1.5, we observe that Flux.1 significantly reduces ”entity omission” errors. For complex prompts like ”a red car and a blue truck”, SD1.5 often merged colors or missed one object, whereas Flux.1 consistently renders both. However, we also note a new failure mode: ”hyper-real hallucination,” where the model generates convincing but irrelevant details. In our pipeline, this manifests as ”inference noise”—if the text does not explicitly contradict these details, the model may over-interpret them, leading to errors in fine-grained sentiment tasks.
@@
In summary, the utility of a generated image is highest when the source text describes concrete, visualizable elements, the T2I model can render these elements with high fidelity, and the downstream task can benefit from this visual grounding. Conversely, the approach is less effective when the text is abstract, the T2I generation is of low quality or semantically misaligned, the textual information is already self-sufficient, or the task itself has low correlation with visual information. This analysis provides a clear answer to RQ6. See Appendix D for additional illustrative case studies. For a detailed discussion of our findings, including their implications and limitations, please see Appendix K.